OpenStudy (anonymous):

A ladder 12 m in length rests against a wall. The foot of the ladder is 3 m from the wall. What is the measure of the angle the ladder forms with the floor?

OpenStudy (raden):



OpenStudy (raden):

use defined of cos to get x cos(x) = 3/12 cos(x) = 1/4 x = arc(cos(1/4)) = ... (use the calculator)